Maria is now 3 years younger that her sister Ann. Four years ago, Maria's age was one-half Ann's age. How old is each girl now?
Answer by ewatrrr(24785)


Hi
Maria is now 3 years younger that her sister Ann.
Let x and (x-3) represent Ann's and Maria's ages NOW respectively
Question states***four years ago...
(x-3) - 4 = x/2
Solving for x
x - 7 = x/2
x/2 = 7
x = 14, Ann's age now. Maria is 11 (14-3)
CHECKING our Answer***
11-4 = 14 = 7